Wesco Incorporated’s only product is a combination fertilizer/weedkiller called GrowNWeed. GrowNWeed is sold nationwide to retail nurseries and garden stores. Zwinger Nursery plans to sell a similar fertilizer/weedkiller compound through its regional nursery chain under its own private label. Zwinger does not have manufacturing facilities of its own, so it has asked Wesco (and several other companies) to submit a bid for manufacturing and delivering a 25,000-pound order of the private brand compound to Zwinger. While the chemical composition of the Zwinger compound differs from that of GrowNWeed, the manufacturing processes are very similar. The Zwinger compound would be produced in 1,000-pound lots. Each lot would require 34 direct labor-hours and the following chemicals: Chemicals Quantity in Pounds AG-5 300 KL-2 220 CW-7 190 DF-6 290 The first three chemicals (AG-5, KL-2, and CW-7) are all used in the production of GrowNWeed. DF-6 was used in another compound that Wesco discontinued several months ago. The supply of DF-6 that Wesco had on hand when the other compound was discontinued was not discarded. Wesco could sell its supply of DF-6 at the prevailing market price less $0.11 per pound selling and handling expenses. Wesco also has on hand a chemical called BH-3, which was manufactured for use in another product that is no longer produced. BH-3, which cannot be used in GrowNWeed, can be substituted for AG-5 on a one-for-one basis without affecting the quality of the Zwinger compound. The BH-3 in inventory has a salvage value of $540. Inventory and cost data for the chemicals that can be used to produce the Zwinger compound are shown below: Raw Material Pounds in Inventory Actual Price per Pound When Purchased Current Market Price per Pound AG-5 26,000 $ 0.71 $ 0.81 KL-2 4,500 $ 0.40 $ 0.45 CW-7 7,800 $ 1.26 $ 1.46 DF-6 5,850 $ 0.55 $ 0.50 BH-3 5,800 $ 0.56 (Salvage) The current direct labor wage rate is $12 per hour. The predetermined overhead rate is based on direct labor-hours (DLH). The predetermined overhead rate for the current year, based on a two-shift capacity with no overtime, is as follows: Variable manufacturing overhead $ 5.00 per DLH Fixed manufacturing overhead 6.80 per DLH Combined predetermined overhead rate $ 11.80 per DLH Wesco’s production manager reports that the present equipment and facilities are adequate to manufacture the Zwinger compound. Therefore, the order would have no effect on total fixed manufacturing overhead costs. However, Wesco is within 120 hours of its two-shift capacity this month. Any additional hours beyond the 120 hours must be done in overtime. If need be, the Zwinger compound could be produced on regular time by shifting a portion of GrowNWeed production to overtime. Wesco’s direct labor wage rate for overtime is $18 per hour. There is no allowance for any overtime premium in the predetermined overhead rate. Required: 1. Wesco has decided to submit a bid for the 25,000 pound order of Zwinger’s new compound. The order must be delivered by the end of the current month. Zwinger has indicated that this is a one-time order that will not be repeated. Calculate the lowest price that Wesco could bid for the order and still exactly cover its incremental manufacturing costs. 2. Refer to the original data. Assume that Zwinger Nursery plans to place regular orders for 25,000-pound lots of the new compound. Wesco expects the demand for GrowNWeed to remain strong. Therefore, the recurring orders from Zwinger would put Wesco over its two-shift capacity. However, production could be scheduled so that 60% of each Zwinger order could be completed during regular hours. As another option, some GrowNWeed production could be shifted temporarily to overtime so that the Zwinger orders could be produced on regular time. Current market prices are the best available estimates of future market prices. Wesco’s standard markup policy for new products is 40% of the full manufacturing cost, including fixed manufacturing overhead. Calculate the price that Wesco, Inc., would quote Zwinger Nursery for each 25,000 pound lot of the new compound, assuming that it is to be treated as a new product and this pricing policy is followed.
Answer #1
1 Lowest price:
$ $
Direct materials:
AG-5 (Note:1) 1917
KL-2 (Note:2) 2475
CW-7 (Note:3) 6935
DF-6 (Note:4) 2827.5
Total direct material cost 14154.5
Direct labor (Note:5) 14580
Variable manufacturing overhead (Note:6) 4250
Total cost (Lowest price for the bid) 32984.5
Note:1
Special order size=25000 pounds
Produced in 1000 pound lots
Number of lots=25000/1000=25 lots
Total quantity required of AG-5=25*300=7500 pounds
A substitute BH-3 available for AG-5 on a one-for-one basis
BH-3 available in inventory=5800 pounds
It's salvage value is relevant cost
Salvage value=$ 540
Quantity of AG-5 to be used=7500-5800=1700 pounds
Cost of AG-5=1700*0.81=$ 1377
Total cost=540+1377=$ 1917
Note:2
Special order size=25000 pounds
Produced in 1000 pound lots
Number of lots=25000/1000=25 lots
Total quantity required of KL-2=25*220=5500 pounds
Cost of KL-2=5500*0.45=$ 2475
Note:3
Special order size=25000 pounds
Produced in 1000 pound lots
Number of lots=25000/1000=25 lots
Total quantity required of CW-7=25*190=4750 pounds
Cost of CW-7=4750*1.46=$ 6935
Note:4
Special order size=25000 pounds
Produced in 1000 pound lots
Number of lots=25000/1000=25 lots
Total quantity required of DF-6=25*290=7250 pounds
Current inventory of DF-6 Can be sold for market price less $ 0.11 per pound
Sale value Is a relevant cost
Sale value=5850*(0.50-0.11)=5850*0.39=$ 2281.5
Balance quantity to be used=7250-5850=1400 pounds
Cost of DF-6=1400*0.39=$ 546
Total relevant cost of DF-6=2281.5+546=$ 2827.5
Note:5
Special order size=25000 pounds
Produced in 1000 pound lots
Number of lots=25000/1000=25 lots
Total labor hours required=25*34=850 hours
Regular hours available=120 hours
Overtime hours=850-120=730 hours
Cost of regular hours=120*12=$ 1440
Cost of overtime hours=730*18=$ 13140
Total direct labor cost=1440+13140=$ 14580
Note:6
Special order size=25000 pounds
Produced in 1000 pound lots
Number of lots=25000/1000=25 lots
Total labor hours required=25*34=850 hours
Variable manufacturing overhead rate=$5 per DLH
Total cost of variable manufacturing overhead=850*5=$ 4250
2 Selling price:
$ $
Direct material:
AG-5 (25*300*0.81) 6075
KL-2 (25*220*0.45) 2475
CW-7 (25*190*1.46) 6935
DF-6 (25*290*0.50) 3625
Total direct material cost 19110
Direct labor:
Regular (25*34*60%*12) 6120
Overtime (25*34*40%*18) 6120
Total direct labor cost 12240
Manufacturing overhead applied (25*34*11.80) 10030
Total manufacturing cost 41380
Add: markup at 40% 16552
Selling price 57932